a demand curve is a straight line sloping downwards at an angle of 45 degrees.we can conclude that its price elasticity is a-1.0 b-0.5 c-constant d-variable
Answer #1

Answer: d - variable

When a demand curve is a straight line sloping downwards at an angle of 45 degrees, the value of its vertical and horizontal intercepts are same.

Tan =Perpendicular / Base

We know that Tan 45o = 1

or, Perpendicular / Base = 1

Perpendicular = Base

In the above graph, price(P) is measured on the vertical axis, and quantity(Q) is measured on the horizontal axis. Let the maximum price is $30. So its corresponding maximum quantity would also be 30 unit. Joining maximum price, and maximum quantity, we get the downward sloping straight line demand curve, 'D1D1', which makes 45o angle to the horizontal axis.

Price elasticity of demand(e) = - [(dQ/Q) / dP/P)]

Or, e = - (dQ/ dP) * (P/Q), [where, dQ/ dP = slope of the inverse demand curve]

Taking modulus of e, i.e., |e| = (dQ/ dP) * (P/Q)

Now, slope of the inverse demand curve = 1/(dP/dQ) = 1 / 1 = 1 [ dP/ dQ = 1]

When price is half of maximum price, i.e., when P = 15, we see its corresponding quantity is also 15.

So at the mid-point on the demand curve, the elasticity, e or |e| = (dQ/ dP) * (P/Q) = 1*1 = 1 = Unitary elasticity

Any point above the mid-point of the demand curve , |e| 1 , as PQ [Price rises, quantity demand decreases.]

Any point below the mid-point of the demand curve , |e| 1 , as PQ [Price decreases, quantity demand rises]

The two extreme points on the demand curve, are point 'T', and point 'L'.

At point T; |e| = ∞(infinity).

At point L; |e| = 0.

The value of price elasticity of demand for normal good is negative. To make the sign positive, we take the modulus of elasticity (e) , i.e., |e|.

________________________________________________________________